OpenStudy (anonymous):

Use the disk or the shell method to find the volume of the solid generated by revolving the region bounded by the graphs of the equations about each given line. y = 6/x^2 y = 0 x = 1 x = 3 What method should I use?

OpenStudy (amistre64):

all of them lol

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

I've tried all of them, and I still get the wrong answer

OpenStudy (anonymous):

The shell method gives me -8pi, and the disk method gives me a ridiculous number

OpenStudy (amistre64):

the question seems to be asking you to do it 4 different times...

OpenStudy (anonymous):

Oh, my question uses the y-axis.

OpenStudy (anonymous):

There's a list of them, and I have already done the one revolving around the x-axis

OpenStudy (amistre64):

is this our region?

OpenStudy (amistre64):

if you wanna spin it round the y axis; shell it. 2pi 6 {S} x(1/x^2) dx ; [1,3]

OpenStudy (anonymous):

Yes...so I thought to use the shell method first....right., that's what I did, and got -8pi

OpenStudy (amistre64):

1/x ints up to ln(x) right? so 12pi ln(x)

OpenStudy (anonymous):

wait---you separated the 6 from the equation, and multiplied by x instead of (x + 1) or some variation of that?

OpenStudy (amistre64):

12pi ln(3) should be it ; since ln(1) = 0

OpenStudy (amistre64):

constants dont get inted; they get pulled out, thats why we dont int pi and 2pi and the like

OpenStudy (amistre64):

we dont derive constants either; so why bother inting them lol

OpenStudy (amistre64):

sine the y axis is alreay x = 0; no need to move anything

OpenStudy (anonymous):

so we don't add or subtract anything from the r(x) factor in [\int\limits_{a}^{b}r(x)h(x)dx\] if y = 0?

OpenStudy (anonymous):

\[2\pi \int\limits_{a}^{b}r(x)h(x)dx\]

OpenStudy (amistre64):

nope; the radius in the shell method is as you move from 1 to 3 in this and the height is 6/x^2

OpenStudy (anonymous):

the radius in the shell method is...? as you move from 1 to 3?

OpenStudy (amistre64):

what are we spinning around?

OpenStudy (anonymous):

the y-axis

OpenStudy (amistre64):

does the y axis the same as the x=0 axis?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

is the y axis that same as the x=0 line?

OpenStudy (anonymous):

y = 0 and x = 0 look the same....oh, okay

OpenStudy (amistre64):

since y axis IS x=0; then what do we move to get this to x=0? nothing right?

OpenStudy (amistre64):

so r(x) = x ; h(x) = 6/x^2 r(x)h(x) = 6/x from 1 to 3

OpenStudy (amistre64):

12pi/x ints to 12pi ln(x) which means the the answer is: 12pi ln(3)

OpenStudy (anonymous):

the answer is right! i get it----what your second to last post was. This makes sense now.

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Thanks! See yah next time